# Projected Augmented Waves (PAW) inspired mixed basis sets for small molecules

### v3

@@ III.2 Slo and SLO basis sets
While four center integrals are very difficult with this basis set three center integrals are numerically manageable [27]. Below we describe practical methods to achieve accurate three center decompositions.
@@ III.2.1 Practical methods for implementation(three center decompositions)
this allows for hybridization between the atomic orbitals such as $sp^{3}$ hybridization common to Carbon (if this effect is not important for the molecule just the diagonal terms suffice). Furthermore for notational reasons we absorb the label $\mu;Elm;E^{\prime}l^{\prime}m^{\prime}$ into $\kappa$. We introduce the projectors $P_{\lambda}$ such that
@@
We now introduce more notation for the single particle basis set where $\left\{\nu\right\}=\left\{Elm\right\}\cup\left\{\mu,l,m,\xi\right\}$ and write the eigenstates are given by
@@ IV Conclusions
In this work - inspired by the fact that many pseudopotential methods are all electron methods in disguise with slowly oscillating basis wave functions and special basis wave functions to account for the environment of the nucleus - we proposed new basis sets for small molecules and crystalline solids. In this basis set we use the LO or lo wave functions introduced recently in [26] (although regular LO or lo basis wave functions will also do [5, 6, 1]) combined either with Gaussians (GTOs and CGFs) and slater basis sets - STO - (for molecules) or plane waves (for solids) to obtain a total basis for the system. This would allow for initial calculations for many molecules as the four center calculation is done for Gaussians and two center density density is done for very generic functions [2, 29, 18] and for slater type orbitals it is possible to introduce appropriate density shape functions. This should help open small molecules for theoretical DFT or HF like calculations and explorations.

### v4

@@ III.2 Slo and SLO basis sets
While four center integrals are very difficult with this basis set three center integrals are numerically manageable [29]. Below we describe practical methods to achieve accurate three center decompositions into shape functions.
@@ III.2.1 Practical methods for implementation (three center decompositions)
this allows for hybridization between the atomic orbitals such as for example the $sp^{3}$ hybridization common to Carbon (if this effect is not important for the molecule just the diagonal terms suffice). Furthermore for notational reasons we absorb the label $\mu;Elm;E^{\prime}l^{\prime}m^{\prime}$ into $\kappa$. We introduce the projectors $P_{\lambda}$ such that
@@
We now introduce more notation for the single particle basis set where $\left\{\nu\right\}=\left\{\mu,Elm\right\}\cup\left\{\mu,l,m,\xi\right\}$ and write the eigenstates are given by
@@ IV Conclusions
In this work - motivated by the fact that many pseudopotential methods are all electron methods in disguise with basis sets of the form of slowly oscillating basis wave functions and special high momentum basis wave functions to account for the environment of the nucleus - we proposed several related new basis sets for small molecules and crystalline solids. In this basis set we use the LO or lo wave functions introduced recently in [28] (although regular LO or lo basis wave functions will also do quite well [5, 6, 1]) combined either with Gaussians (GTOs and CGFs) or Slater basis sets - STO - (for molecules) or plane waves (for solids) to obtain a total basis for the system. This would allow for initial calculations for many molecules as the four center calculation is done for Gaussians and two center integrals have been done for very generic functions [2, 31, 18] and for Slater type orbitals it is possible to introduce appropriate density shape functions (which we extended). This should help open small molecules for theoretical DFT or HF like calculations and explorations.
Appendix A Gaussian basis sets (review)
We are interested in four center integrals for the Coulomb problem [20, 18]:
$I_{\alpha\beta\gamma\delta}=$
$\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\frac{1}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}\varphi_{\alpha}^{*}\left(\mathbf{r}_{1}\right)\varphi_{\beta}\left(\mathbf{r}_{1}\right)\varphi_{\gamma}^{*}\left(\mathbf{r}_{2}\right)\varphi_{\delta}\left(\mathbf{r}_{2}\right)$ (28)
This problem scales as $N^{4}$ (where $N$ is the size of the basis set) and must be evaluated for non-crystalline materials (molecules) where the Weinert method does not directly apply. As such it is often the bottleneck for many calculations [20, 18]. We can introduce spherical Gaussians to simplify these integrals:
$G_{lm,n,\mathbf{R},p}\left(\mathbf{r}\right)=Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{R}}\right)\exp\left(-p\left(\mathbf{r}-\mathbf{R}\right)^{2}\right)\left|\mathbf{r}-\mathbf{R}\right|^{2n+l}$ (29)
We can also introduce cartesian Gaussians:
$G_{a,b,c,\mathbf{R},p}\left(\mathbf{r}\right)$
$=\exp\left(-p\left(\mathbf{r}-\mathbf{R}\right)^{2}\right)\left(x-\mathbf{R}_{x}\right)^{a}\left(y-\mathbf{R}_{y}\right)^{b}\left(z-\mathbf{R}_{z}\right)^{c}$ (30)
We now have that:
$G_{lm,n,\mathbf{R},p}\left(\mathbf{r}\right)=\sum_{a+b+c=l+2n}A_{abc}^{lmn}G_{a,b,c,\mathbf{R},p}\left(\mathbf{r}\right),$ (31)
so we might as well focus on cartesian Gaussians for the wave functions in the integral in Eq. (28). Now we have that:
$G_{a,b,c,\mathbf{R},p}\left(\mathbf{r}\right)G_{a^{\prime},b^{\prime},c^{\prime},\mathbf{R}^{\prime},p^{\prime}}\left(\mathbf{r}\right)$
$=\sum_{ABC}P_{ABC}^{abc,\mathbf{R},\mathbf{R}^{\prime}}G_{ABC\bar{\mathbf{R}},p+p^{\prime}}\left(\mathbf{r}\right)$ (32)
Where
$\bar{\mathbf{R}}=\frac{p\mathbf{R}+p^{\prime}\mathbf{R}^{\prime}}{p+p^{\prime}}$ (33)
As such for cartesian Gaussians, up to linear sums, we may as well focus on two center integrals of the form:
$I_{\alpha\beta\gamma\delta}=$
$=\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\frac{1}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}G_{abc,\mathbf{R},p}\left(\mathbf{r}_{1}\right)G_{a^{\prime}b^{\prime}c^{\prime},\mathbf{R}^{\prime},p^{\prime}}\left(\mathbf{r}_{2}\right)$
$=\sum\kappa_{abc}^{ABC}\frac{\partial^{A+B+C}}{\partial\mathbf{R}_{x}^{A}\partial\mathbf{R}_{y}^{B}\partial\mathbf{R}_{z}^{C}}\kappa_{a^{\prime}b^{\prime}c^{\prime}}^{A^{\prime}B^{\prime}C^{\prime}}\frac{\partial^{A^{\prime}+B^{\prime}+C^{\prime}}}{\partial\mathbf{R}^{\prime}{}_{x}^{A^{\prime}}\partial\mathbf{R}^{\prime}{}_{y}^{B^{\prime}}\partial\mathbf{R}^{\prime}{}_{z}^{C^{\prime}}}\times$
×$\times\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\frac{1}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}G_{000,\mathbf{R},p}\left(\mathbf{r}_{1}\right)G_{000,\mathbf{R}^{\prime},p^{\prime}}\left(\mathbf{r}_{2}\right)$ (34)
We now introduce the Boys’ function:
[…]
Then we have that:
$\int d^{3}\mathbf{r}_{1}\int d^{3}\mathbf{r}_{2}\frac{1}{\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|}G_{000,\mathbf{R},p}\left(\mathbf{r}_{1}\right)G_{000,\mathbf{R}^{\prime},p^{\prime}}\left(\mathbf{r}_{2}\right)$
$=\sqrt{\frac{4\alpha}{\pi}}F_{0}\left(\alpha\left|\mathbf{R}-\mathbf{R}^{\prime}\right|^{2}\right)$ (36)
With:
$\alpha=\frac{p\cdot p^{\prime}}{p+p^{\prime}}$ (37)
As such four center integrals may be efficiently evaluated with Gaussian wave functions.
?refname?
  • 1 [2000] E. Sjostedt, L. Nordstrom, and D. J. Singh, Sol. Sta. Comm. 114, 15 (2000).
  • 2 [2019] S. M. Blinder and D. E. Huse, Mathematical Physics in Theoretical Chemistry, (Elsevier Inc. Amsterdam, 2019).
  • 3 [2020] R. M. Martin, Electronic Structures Basic Theory and Practical Methods (Cambridge University Press, Cambridge, 2020).
  • 4 [2009] D. Marx and J. Hutter, Ab Initio Molecular Dynamics Basic Theory and Advanced Methods (Cambridge University Press, Cambridge, 2009).
  • 5 [1991] D. J. Singh, Phys. Rev. B 43, 6388 (1991).
  • 6 [2006] D. J. Singh and D. Nordstrom, Planewaves, pseudopotentials, and the LAPW method (Springer, New York, 2006).
  • 7 [2014] G. Michalicek, Extending the precision and efficiency of all-electron full-potential linearized augment plane-wave density functional theory (Aachen University, 2014, thesis).
  • 8 [2013] G. Michalicek, M. Betzinger, C. Freidrich and S. Blugel, Comp. Phys. Comm. 184, 2670 (2013).
  • 9 [1975] O. K. Andersen, Phys. Rev. B 12, 3060 (1975).
  • 10 [1984] H. L. Skriver, The LMTO method Muffin-Tin Orbitals and Electronic Structure (Springer, New York, 1984).
  • 11 [1984] O. K. Andersen, and O. Jepsen, Phys. Rev. Lett. 53, 2571 (1984).
  • 12 [2003] O. K. Andersen, T. S.-Dasgupta, and S. Ezhof, Bull. Mat. Sci. 26, 19 (2003).
  • 13 [1954] W. Kohn, N. Rostoker, Phys. Rev. 94, $1$ (1954).
  • 14 [1947] J. Korringa, Physica 13, 392 (1947).
  • 15 [1989] J. M. Soler, and A. R. Williams, Phys. Rev. B 40, 1560 (1989).
  • 16 [1990] J. M. Soler and A. R. Williams, Phys. Rev. B 42, 9728 (1990).
  • 17 [1967] T. L. Loucks, Augmented Plane Wave Method (W. A. Benjamin Inc., New York, 1967).
  • 18 [1975] T. Helgaker, P. Jorgensen, J. Olsen, Molecular Electronic Structure Theory (John Wiley & Sons LTD. New York, 2000).
  • 19 [2014] I. N. Levine, Quantum Chemistry (Pearson Education Inc., Boston, 2014).
  • 20 [1995] T. Helgakar and P. R. Taylor, “Gaussian Basis sets and molecular integrals” in Modern Electronic structure theory part II, D. R. Yarkony (eds.) (World Scientific, Singapore, 1995).
  • 21 [1990] D. Feller and E. R. Davidson, Rev. Comp. Chem. 1, 1 (1990).
  • 22 [1972] W. J. Hehre, R. Ditchfield, J. A. Pople, J. Chem. Phys. 56, 2257 (1972).
  • 23 [1973] P. C. Hariharan and J. A. Pople, Theor. Chem. Acta. 28, 213 (1973).
  • 24 [1994] P. E. Blochl, Phys. Rev. B 50, 17953 (1994).
  • 25 [1996] G. Kresse, J. Furthmuller, Phys. Rev. B 54, 11169 (1996).
  • 26 [2024] G. Goldstein, arXiv 2405.11926
  • 27 [2024] G. Goldstein, arXiv 2407.15299
  • 28 [2024] G. Goldstein, arXiv 2403.15954.
  • 29 [2001] W. Koch, and M. C. Holthausen, A chemist’s guide to density functional theory (Wiley, new York, 2001).
  • 30 [1950] S. F. Boys, Proc. Roy. Soc. A200, 542 (1950).
  • 31 [2024] G. Goldstein, arXiv 2404.08493